• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kelly

    Log entry time 16:01:36 on June23,2000

    Entry number 44726

    keyword=day shift summary





    production running at thcm = -50

    1000: begin laser spot move
    1135: MCC reported that spot move was accomplished by about 1050
    and went well; however, they are having problems with the power
    supply for a trim magnet. The same problem occurred yesterday.
    No estimate yet for the time needed to track down and repair the
    problem once and for all.
    1210: beam back. Request and get about 80 uA!
    1300: unstable beam, frequent trips
    1330: MCC reports injector BCM trips. Stop to investigate.
    Only a few spurts of beam during rest of this shift. Experts called in.

    run events stop comments
    2471 1.3 M 0813
    2472 1.3 M 0905
    2473 1.3 M 1001
    2474 1.3 M 1253 new spot, 83 uA
    2475 0.6 M 1330 very instable beam for entire run
    2476 0.4 M 1521 ended run for accelerator studies